# MMO-IG: Multi-Class and Multi-Scale Object Image Generation for Remote Sensing

### v1

@@ Abstract
The rapid advancement of deep generative models (DGMs) has significantly advanced research in computer vision, providing a cost-effective alternative to acquiring vast quantities of expensive imagery. However, existing methods predominantly focus on synthesizing remote sensing (RS) images aligned with real images in a global layout view, which limits their applicability in RS image object detection (RSIOD) research. To address these challenges, we propose a multi-class and multi-scale object image generator based on DGMs, termed MMO-IG, designed to generate RS images with supervised object labels from global and local aspects simultaneously. Specifically, from the local view, MMO-IG encodes various RS instances using an iso-spacing instance map (ISIM). During the generation process, it decodes each instance region with iso-spacing value in ISIM—corresponding to both background and foreground instances—to produce RS images through the denoising process of diffusion models. Considering the complex interdependencies among MMOs, we construct a spatial-cross dependency knowledge graph (SCDKG). This ensures a realistic and reliable multidirectional distribution among MMOs for region embedding, thereby reducing the discrepancy between source and target domains. Besides, we propose a structured object distribution instruction (SODI) to guide the generation of synthesized RS image content from a global aspect with SCDKG-based ISIM together. Extensive experimental results demonstrate that our MMO-IG exhibits superior generation capabilities for RS images with dense MMO-supervised labels, and RS detectors pre-trained with MMO-IG show excellent performance on real-world datasets.
@@
Figure 1: Illustration of the generation of RS images containing MMOs by the proposed MMO-IG. Notably, each RS object is modeled by a unique $\mathcal{P}_{\rm scg}$ according to the corresponding realistic geometric characteristics.
@@ I Introduction
Object detection in remote sensing (RS) images [1, 2, 3, 4] is an important task in earth observation technology, providing essential support for various application scenarios, such as disaster monitoring, military reconnaissance, traffic monitoring, and smart cities. This task aims to determine object categories and locations based on input optical RS images. With the progress of deep learning-based generic detection technology [5, 6, 7], remote sensing image object detection (RSIOD) research has achieved significant development, where analyzing complex backgrounds and focusing on varied class and scale foreground objects simultaneously is becoming achievable as long as the algorithm that is well-trained based on sufficient and high-quality data. However, the high cost of acquiring satellite RS images and the labor-intensive nature of image annotation limit the availability of data for this research area, making it challenging to adequately train algorithms. This widespread data limitation issue in the field of RS has consistently hindered researchers from making further breakthroughs.
@@
Recently, the rapidly developed deep generative models (DGMs) [8, 9, 10] provide a solution for alleviating the above issue. They aim to generate realistic and high-quality images. It enhances the accessibility of data in resource-constrained environments significantly, which makes it possible to build RS generative models (RSGMs) [11, 12] for alleviating the data limitation of RSIOD. While existing RSGMs follow generic DGMs to focus on the achievement of text-to-image or layout-to-image [13, 14, 15, 16], where the former generates RS images according to text descriptions prompt and the latter is designed to render RS images based on a given layout image prompt (e.g., depth map, sketch, road, edge, etc.). It means they are hard to provide qualified data with instance-level labels that can model complex interdependencies among different RS objects for the RSIOD task.
@@
Following the above issue, we introduce an instance-aware image generator for providing sufficient and high-quality data with multi-class and multi-scale object (MMO) supervision information for RS detection, namely MMO-IG. Specifically, inspired by popular diffusion-based controlled generative models [10, 17, 18, 19], we formulate the MMO-IG as a controllable generation problem. Different from previous methods that were designed to generate RS images whose content adheres to global layout requirements, the proposed MMO-IG is demanded to focus on each object instance’s class, location, and scale in a local view while considering the interdependencies among different instances from a global aspect. Considering that the existing text and layout prompt cannot represent MMOs, an iso-spacing instance map (ISIM) is introduced as the control signal for guidance in our model in generating RS images containing MMPs, where instances with different classes are encoded via various iso-spacing valued grayscale regions. During testing, MMO-IG decodes these regions into instances with different classes according to regions’ grayscale values while keeping the location and scale characteristics of the corresponding regions on generated instances. Besides, distinct from previous RSGMs concentrate on pursuing a correct corresponding between the given text or layout prompt and the generated RS image, there exist complex interdependencies among different objects with the same class and also in different classes of objects, which demands our model to ensure a rational interdependent distribution among the MMOs on the generated RS image. Based on this consideration, we construct a spatial-cross dependency knowledge graph (SCDKG) to formulate the complex interdependencies among different RS objects. During testing, we synthesize an ISIM with rational interdependent distribution among the MMOs under the guidance of SCDKG to render the RS image via MMO-IG more accurately reflective of reality. Based on the above proposed ISIM and SCDKG, we construct MMO-IG for generating mass RS images with dense instance-level labels, which provides a solution for alleviating the data limitation problem that exists in the RSIOD field. Furthermore, the hallucination problem in deep generative models (DGMs) can cause remote sensing (RS) objects to appear incorrectly in the background, leading to discrepancies between instance-level labels (i.e., ISIM) and the generated RS images, which negatively impacts downstream tasks. To address this issue, SODI is introduced to guide the image generation process by considering the global perspective of image style and instance characteristics, ensuring both accuracy and control.
@@ II-A Remote Sensing Image Object Detection
Object detection is a key topic in computer vision, which provides sufficient support to analyze image context. Inspired by generic object detection approaches [20, 5, 6, 7, 21, 22, 23], researchers designed RSIOD frameworks according to the specific geometric characteristics of RS objects, which become critical techniques for RS scenario applications. Existing RSIOD methods can be classified into small [24, 3, 25] and oriented [2, 26, 27] object detection methods roughly according to the object characteristics. The former concentrated on designing the feature fusion structure [28, 29, 30] or introducing attention mechanism into frameworks [31, 32, 33] to obtain strong expression ability to help the model distinguish small targets from complex backgrounds more effectively. The latter primarily aimed to represent oriented object boundaries [34, 35] accurately to avoid superfluous background interference. Besides, some works [36, 37, 38] focus on optimizing the whole detection framework to achieve performance balance between accuracy and efficiency. Although these methods achieve competitive performance on multiple public datasets, the limited availability of data in many scenarios hinders their ability to maintain superior performance.
@@ II-B Deep Generative Models
Figure 2: Overall pipeline of MMO-IG for generating RS images with dense instance-level bounding box labels. It first synthesizes rational spatial geometric characteristics of MMOs via SCDKG. They then are encoded via the designed ISIM while describing the RS image content through SODI. In the end, following the diffusion model to decode the ISIM to RS image contained MMOs under the guidance of SODI. Notably, each RS object is modeled by a unique $\mathcal{P}_{\rm scg}$ according to the corresponding realistic geometric characteristics.
@@ III-A Overall Pipeline
Figure 3: Illustration of the proposed SCDKG, which models complex interdependencies among objects of different classes via $\rm p_{id}$ matrix and their diverse spatial geometric characteristics via $\mathcal{P}_{\rm sgc}$. Notably, each RS object is modeled by a unique $\mathcal{P}_{\rm scg}$ according to the corresponding realistic geometric characteristics.
@@ III-B Spatial-Cross Dependency Knowledge Graph
Upon determining the initial object class, SCDKG proceeds to define object attributes (including aspect ratio, scale, and location) by adhering to the spatial geometric characteristics of realistic RS objects and the initial class. Concretely, it models each class object through the spatial geometric characteristics probability density function $\mathcal{P}_{\rm scg}$. For attributes like aspect ratio and scale, SCDKG employs a one-dimensional probability density function after analyzing the relevant geometric features of all instances in the dataset. The location attribute is modeled by decomposing the instance center point coordinates into two dimensions, $x$ and $y$, and fitting them with a two-dimensional probability density function. Utilizing $\mathcal{P}^{\rm all}_{\rm scg}$ for all RS objects, SCDKG assigns attributes according to the class-specific $\mathcal{P}_{\rm scg}$.
@@
The $\rm p_{id}$ values for different object class pairs form an interdependency probability matrix (illustrated in Fig.2 $\rm p_{id}$ matrix). Based on the previous instance class, SCDKG samples the next instance class from this matrix and assigns attributes through the corresponding $\mathcal{P}_{\rm scg}$. Finally, the model determines a rational instance density (i.e., ’N’ in Step 1 of Fig. 2) using the probability density function $\mathcal{P}_{\rm in}$.
@@ III-C Iso-Spacing Instance Map
Given a remote sensing (RS) dataset containing $\rm M$ classes, the ISIM method initially assigns each class a unique numerical identifier ranging from $\left[0\sim\rm{M}\right]$. This identifier is then used in an arithmetic interval grayscale assignment strategy. The grayscale value corresponding to each class can be computed as the following equation:
where $\rm m$ represents the unique numerical identifier for each class instance and $\rm v_{gray}$ is the corresponding grayscale value. With the determined class encoding for all kinds of instances in a specific dataset, MMO attributes from Step 1 in Fig. 2 are utilized to synthesize ISIM. Concretely, ISIM encodes each object based on regions corresponding to the class-specific grayscale values and spatial geometric characteristics (i.e., aspect ratio, scale, and location) on a grayscale image.
@@ III-E Decoding ISIM to Generate RS Image
Diffusion models, a popular class of deep generative models (DGMs) [45, 8, 9, 10], reconstruct images via a denoising process applied to noisy inputs. These models offer advantages such as stable training, high-quality sample generation, tractable likelihood estimation, versatility, and robustness to hyperparameter settings, outperforming VAE [39] and GAN [42] models. Consequently, MMO-IG adopts the stable diffusion [9] framework for constructing its generative network.
As illustrated in Step 3 of Fig. 2, the generative component primarily comprises two pairs of residual encoders and decoders. Each encoder and decoder contains 12 stable diffusion layers, interconnected by a central layer consisting of a single stable diffusion layer. The generative blocks are initialized with pre-trained weights from stable diffusion [9]. The trainable generative blocks replicate the structure of the fixed blocks, with initial weights set by copying from the fixed block and zeroing, respectively.
During the generation process, the network takes ISIM, SODI, and $z^{T}$ as inputs to produce RS images. ISIM and SODI are embedded into feature vectors via a lightweight network $\varepsilon$ and CLIP [49], respectively. The network $\varepsilon$ consists of four convolutional layers with 4$\times$4 kernels and 2$\times$2 strides $\rm Conv^{16}_{4,2},Conv^{32}_{4,2},Conv^{64}_{4,2},andConv^{128}_{4,2}$. The $z^{T}$ represents the latent feature of the noisy image, efficiently obtained by applying the trained $\varepsilon$ on the noisy image.
After $T$-1 iterative denoising steps through the residual blocks, MMO-IG decodes each region in ISIM into the corresponding object using the decoded image latent feature $z$. The final RS image is reconstructed by applying $D$ (a convolution block proposed in [50] that is used for transferring small-sized features into large RGB image) to $z$.
@@ IV-B Ablation Study
Effectiveness of SODI. Considering the hallucination problem in deep generative models (DGMs) can cause remote sensing (RS) objects to be inaccurately represented in the background, resulting in inconsistencies between instance-level labels (ISIM) and the generated RS images (as shown in Fig. 6). This discrepancy adversely affects downstream tasks. Therefore, SODI is implemented to guide the image generation process by incorporating a global perspective on image style and instance characteristics, thereby ensuring precision and control. To verify the effectiveness of SODI in ensuring the alignment between ISIM and the generated RS images, we show the testing quantitative metrics of our model with different text conditions in Table I, where ‘NULL’ and ‘BLIP’ denote the text condition is null and an image description generated by BLIP [52]. It can be found that MMO-IG with ‘NULL’ text condition makes it hard to control the accuracy of object class and number on the generated RS images, where the $\rm Acc_{c}$ and $\rm Acc_{n}$ are 92.8 and 93.9 respectively. It means that ISIM, the instance-level label, enjoys low-quality correspondence with the generated RS images, further influencing the following downstream task.
@@
SCDKG consists of $\rm p_{id}$ matrix that models interdependencies among all objects of different classes (the $\rm p_{id}$ matrix of DIOR dataset is shown in Fig. 7) and $\mathcal{P}_{\rm scg}$ that formulates diverse spatial geometric characteristics of each object class individually. Notably, $\mathcal{P}_{\rm scg}$ models the aspect ratio, scale, and location characteristics of RS objects separately. It effectively avoids influence brought by irrational characteristics (e.g., aspect ratio and positional distribution that do not conform to real-world object spatial geometric characteristics visualized in Fig. 8). To verify the effectiveness of SCDKG, we test the performance enhancement brought by each factor in SCDKG and visualize the results in Table II. For $\mathcal{P}_{\rm scg}$, It can be found that all of the factors can bring performance gains for the generated RS images. Specifically, the modeling of the complex interdependencies among different RS objects makes a more rational object distribution, which helps our model achieve 3.8 and 1.1 improvements in $\rm FID_{zs}$ and CAS on DIOR dataset and 6.8 and 1.0 gains in $\rm FID_{zs}$ and CAS on DIOR-R dataset. Notably, $\rm FID_{zs}$ and $\rm CAS_{zs}$ denote the zero-shot FID and zero-shot CAS, which is introduced to evaluate the quality of the generated images that not in the test set. The improvements brought by $\rm p_{id}$ matrix demonstrate the effectiveness of the interdependencies modeling.
@@
Besides, the rational characteristics of the object aspect ratio, scale, and location bring 8.8, 5.8, and 14.0 improvements in $\rm FID_{zs}$ and 1.4, 3.1, and 5.6 gains in $\rm CAS_{zs}$. This is mainly because the synthesized images enjoy more rational and realistic object spatial characteristics than real RS images, which makes it easier for our model to learn the feature distribution and bring significant performance improvements. Meanwhile, the combination of the three characteristics also can bring further improvements in both $\rm FID_{zs}$ and $\rm CAS{zs}$. Specifically, the model equipped with SCDKG gains 26.7 and 17.6 in $\rm FID_{zs}$ and $\rm CAS{zs}$ on DIOR dataset and 29.4 and 19.5 in $\rm FID_{zs}$ and $\rm CAS{zs}$ on DIOR-R dataset. Significant performance improvements verify the effectiveness of the proposed SCDKG in the generation task of RS images containing MMOs. Some qualified samples are shown in Fig. 9, it can be found that SCDKG enables ensuring rational spatial geometric characteristics for objects with different classes to enhance the reality of generated images.

### v3

@@ Abstract
The rapid advancement of deep generative models (DGMs) has significantly advanced research in computer vision, providing a cost-effective alternative to acquiring vast quantities of expensive imagery. However, existing methods predominantly focus on synthesizing remote sensing (RS) images aligned with real images in a global layout view, which limits their applicability in RS image object detection (RSIOD) research. To address these challenges, we propose a multi-class and multi-scale object image generator based on DGMs, termed MMO-IG, designed to generate RS images with supervised object labels from global and local aspects simultaneously. Specifically, from the local view, MMO-IG encodes various RS instances using an iso-spacing instance map (ISIM). During the generation process, it decodes each instance region with iso-spacing value in ISIM—corresponding to both background and foreground instances—to produce RS images through the denoising process of diffusion models. Considering the complex interdependencies among MMOs, we construct a spatial-cross dependency knowledge graph (SCDKG). This ensures a realistic and reliable multidirectional distribution among MMOs for region embedding, thereby reducing the discrepancy between source and target domains. Besides, we propose a structured object distribution instruction (SODI) to guide the generation of synthesized RS image content from a global aspect with SCDKG-based ISIM together. Extensive experimental results demonstrate that our MMO-IG exhibits superior generation capabilities for RS images with dense MMO-supervised labels, and RS detectors pre-trained with MMO-IG show excellent performance on real-world datasets. Code is available at https://github.com/omtcyang/MMO-IG.
@@
Figure 1: Illustration of the generation of RS images containing MMOs by the proposed MMO-IG. Notably, each RS object is modeled by a unique $\mathcal{P}_{\rm sgc}$ according to the corresponding realistic geometric characteristics.
@@ I Introduction
Object detection in remote sensing (RS) images [1, 2, 3, 4] is an important task in earth observation technology, providing essential support for various application scenarios, such as disaster monitoring, military reconnaissance, traffic monitoring, and smart cities. This task aims to determine object categories and locations based on input optical RS images. With the progress of deep learning-based generic detection technology [5, 6, 7, 8], remote sensing image object detection (RSIOD) research has achieved significant development, where analyzing complex backgrounds and focusing on varied class and scale foreground objects simultaneously is becoming achievable as long as the algorithm that is well-trained based on sufficient and high-quality data. However, the high cost of acquiring satellite RS images and the labor-intensive nature of image annotation limit the availability of data for this research area, making it challenging to adequately train algorithms. This widespread data limitation issue in the field of RS has consistently hindered researchers from making further breakthroughs.
@@
Recently, the rapidly developed deep generative models (DGMs) [9, 10, 11] provide a solution for alleviating the above issue. They aim to generate realistic and high-quality images. It enhances the accessibility of data in resource-constrained environments significantly, which makes it possible to build RS generative models (RSGMs) [12, 13] for alleviating the data limitation of RSIOD. While existing RSGMs follow generic DGMs to focus on the achievement of text-to-image or layout-to-image [14, 15, 16, 17], where the former generates RS images according to text descriptions prompt and the latter is designed to render RS images based on a given layout image prompt (e.g., depth map, sketch, road, edge, etc.). Traditional approaches relying on text or layout prompts face three fundamental limitations when handling Multi-Class and Multi-Scale Objects (MMOs): (1) Layout prompts struggle to encode instance-specific scale variations and precise spatial relationships, often producing rigid object arrangements inconsistent with real RS scenes; (2) Existing methods prioritize global prompt-image alignment while neglecting local interdependencies between adjacent objects; (3) Fixed spatial constraints in layout guidance limit flexible adaptation to diverse RS observation geometries. It means they are hard to provide qualified data with instance-level labels that can model complex interdependencies among different RS objects for the RSIOD task.
@@
Following the above issue, we introduce an instance-aware image generator for providing sufficient and high-quality data with multi-class and multi-scale object (MMO) supervision information for RS detection, namely MMO-IG. To overcome the limitations of traditional guidance signals, we propose the Iso-spacing Instance Map (ISIM) – a novel control mechanism where distinct grayscale regions with unique iso-spacing values encode different object classes. This approach preserves three critical instance attributes: (1) Spatial coordinates through region centroids, (2) Object scales via region areas, and (3) Class identities by grayscale intensities, enabling simultaneous control of multiple object instances while maintaining flexible placement. Specifically, inspired by popular diffusion-based controlled generative models [11, 18, 19, 20], we formulate the MMO-IG as a controllable generation problem. Different from previous methods that were designed to generate RS images whose content adheres to global layout requirements, the proposed MMO-IG is demanded to focus on each object instance’s class, location, and scale in a local view while considering the interdependencies among different instances from a global aspect. Considering that the existing text and layout prompt cannot represent MMOs, an iso-spacing instance map (ISIM) is introduced as the control signal for guidance in our model in generating RS images containing MMPs, where instances with different classes are encoded via various iso-spacing valued grayscale regions. During testing, MMO-IG decodes these regions into instances with different classes according to regions’ grayscale values while keeping the location and scale characteristics of the corresponding regions on generated instances.
Existing RSGMs predominantly focus on pairwise object relationships, failing to capture the complex network of dependencies in real RS scenes where objects form hierarchical clusters (e.g., vehicle groups around buildings) and exhibit cross-class interactions (e.g., ships near ports). To address this, we develop the Spatial-Cross Dependency Knowledge Graph (SCDKG) that explicitly models four types of spatial relationships: (1) Intra-class proximity constraints, (2) Cross-class attraction/repulsion forces, (3) Orientation-aware co-occurrence patterns, and (4) Scale compatibility rules.Besides, distinct from previous RSGMs concentrate on pursuing a correct corresponding between the given text or layout prompt and the generated RS image, there exist complex interdependencies among different objects with the same class and also in different classes of objects, which demands our model to ensure a rational interdependent distribution among the MMOs on the generated RS image. Based on this consideration, we construct a spatial-cross dependency knowledge graph (SCDKG) to formulate the complex interdependencies among different RS objects. During testing, we synthesize an ISIM with rational interdependent distribution among the MMOs under the guidance of SCDKG to render the RS image via MMO-IG more accurately reflective of reality. Based on the above proposed ISIM and SCDKG, we construct MMO-IG for generating mass RS images with dense instance-level labels, which provides a solution for alleviating the data limitation problem that exists in the RSIOD field. Furthermore, the hallucination problem in deep generative models (DGMs) can cause remote sensing (RS) objects to appear incorrectly in the background, leading to discrepancies between instance-level labels (i.e., ISIM) and the generated RS images, which negatively impacts downstream tasks. To address this issue, SODI is introduced to guide the image generation process by considering the global perspective of image style and instance characteristics, ensuring both accuracy and control.
@@ II-A Remote Sensing Image Object Detection
Object detection is a key topic in computer vision, which provides sufficient support to analyze image context. Inspired by generic object detection approaches [21, 22, 23, 24, 25, 26], researchers designed RSIOD frameworks according to the specific geometric characteristics of RS objects, which become critical techniques for RS scenario applications. Existing RSIOD methods can be classified into small [27, 3, 28] and oriented [2, 29, 30] object detection methods roughly according to the object characteristics. The former concentrated on designing the feature fusion structure [31, 32, 33] or introducing attention mechanism into frameworks [34, 35, 36] to obtain strong expression ability to help the model distinguish small targets from complex backgrounds more effectively. The latter primarily aimed to represent oriented object boundaries [37, 38] accurately to avoid superfluous background interference. Besides, some works [39, 40, 41] focus on optimizing the whole detection framework to achieve performance balance between accuracy and efficiency. Although these methods achieve competitive performance on multiple public datasets, the limited availability of data in many scenarios hinders their ability to maintain superior performance.
@@ II-B Deep Generative Models
Figure 2: Overall pipeline of MMO-IG for generating RS images with dense instance-level bounding box labels. It first synthesizes rational spatial geometric characteristics of MMOs via SCDKG. They then are encoded via the designed ISIM while describing the RS image content through SODI. In the end, following the diffusion model to decode the ISIM to RS image contained MMOs under the guidance of SODI. Notably, each RS object is modeled by a unique $\mathcal{P}_{\rm sgc}$ according to the corresponding realistic geometric characteristics.
@@ III-A Overall Pipeline
Figure 3: Illustration of the proposed SCDKG, which models complex interdependencies among objects of different classes via $\rm p_{id}$ matrix and their diverse spatial geometric characteristics via $\mathcal{P}_{\rm sgc}$. Notably, each RS object is modeled by a unique $\mathcal{P}_{\rm sgc}$ according to the corresponding realistic geometric characteristics.
@@ III-B Spatial-Cross Dependency Knowledge Graph
Upon determining the initial object class, SCDKG proceeds to define object attributes (including aspect ratio, scale, and location) by adhering to the spatial geometric characteristics of realistic RS objects and the initial class. The spatial geometric characteristics are modeled through three core probability functions: $\mathcal{P}_{\rm sgc}$ (Spatial Geometric Characteristics) describes instance attributes through location, scale, and aspect ratio distributions. Specifically, the location function fits center point coordinates, the scale function models object lengths, and the length-to-width ratio function determines size proportions. Meanwhile, $\mathcal{P}_{\rm ic}$ (Probability of Instance Class) governs class occurrence likelihoods, $\mathcal{P}_{\rm in}$ (Probability of Instance Number) regulates per-class instance counts, and $\rm p_{id}$ (Probability of Interdependency) quantifies class co-occurrence relationships.
Concretely, it models each class object through the spatial geometric characteristics probability density function $\mathcal{P}_{\rm sgc}$. For attributes like aspect ratio and scale, SCDKG employs a one-dimensional probability density function after analyzing the relevant geometric features of all instances in the dataset. The location attribute is modeled by decomposing the instance center point coordinates into two dimensions, $x$ and $y$, and fitting them with a two-dimensional probability density function. Utilizing $\mathcal{P}^{\rm all}_{\rm sgc}$ for all RS objects, SCDKG assigns attributes according to the class-specific $\mathcal{P}_{\rm sgc}$.
@@
The $\rm p_{id}$ values for different object class pairs form an interdependency probability matrix (illustrated in Fig.2 $\rm p_{id}$ matrix). Based on the previous instance class, SCDKG samples the next instance class from this matrix and assigns attributes through the corresponding $\mathcal{P}_{\rm sgc}$. Finally, the model determines a rational instance density (i.e., ’N’ in Step 1 of Fig. 2) using the probability density function $\mathcal{P}_{\rm in}$.
@@ III-C Iso-Spacing Instance Map
Given a remote sensing (RS) dataset containing $M$ classes, the ISIM method initially assigns each class a unique numerical identifier ranging from $\left[0\sim{M}\right]$. This identifier is then used in an arithmetic interval grayscale assignment strategy. The grayscale value corresponding to each class can be computed as the following equation:
$v_{gray}(m)=\left\lfloor\frac{255}{M}m\right\rfloor,$ (1)
[…]
@@ III-E Decoding ISIM to Generate RS Image
The popular DGMs, diffusion models [48, 9, 10, 11], reconstruct images via a denoising process on noisy counterparts. Considering the diffusion models’ advantages of stable training, high-quality sample generation, tractable likelihood estimation, versatility, and robustness to hyperparameter settings compared with VAE [42] and GAN [45] models, MMO-IG follows the stable diffusion [10] structure to construct the corresponding generative network. As shown in Step 3 of Fig. 2, the generative part consists of two couples of res encoder and res decoder mainly.
In designing our residual encoder (res encoder) and residual decoder (res decoder), we drew inspiration from the architecture of the stable diffusion model. Specifically, the res encoder and res decoder each contain 12 blocks, with the entire model comprising 25 blocks, including the middle block. Among these 25 blocks, 8 blocks are convolutional layers responsible for downsampling or upsampling, while the remaining 17 blocks are the main functional blocks. Each main block consists of 4 residual network layers and 2 Vision Transformers (ViTs). Notably, each ViT incorporates multiple cross-attention and self-attention mechanisms, which further enhance the model’s expressive power and flexibility.
The encoder and decoder are composed of 12 stable diffusion layers respectively and they are connected by a middle layer that of 1 stable diffusion layer. The frozen generative blocks are initialized through the pre-trained weights from stable diffusion [10]. The structures of trainable generative blocks are copied from the right ones and the two block weights are initialized by coping the right block and zero, respectively.
In the generation process, the generative network takes ISIM, SODI, and $z^{T}$ as input to achieve the RS images, where ISIM and SODI are embedding into feature vectors via a lightweight network $\varepsilon$ and CLIP [52] respectively, where $\varepsilon$ consists of four convolution layers with 4$\times$4 kernels and 2$\times$2 strides $\rm Conv^{16}_{4,2},Conv^{32}_{4,2},Conv^{64}_{4,2},andConv^{128}_{4,2}$. $z^{T}$ is the feature vector of noisy image that can be efficiently obtained by conducting the well-trained $\varepsilon$ on the noisy image. After T-1 times iterative denoising through res blocks, MMO-IG decoding each region in ISIM into the corresponding object with the decoded image feature vector $z$, and the final RS image reconstructed by conducting $D$ (that proposed in  [53]) on $z$.
Given the clean input image $z_{0}$, diffusion models gradually corrupt the image through $t$ successive noising steps, producing the noisy latent $z_{t}$ where $t\in\{1,...,T\}$ denotes the timestep index. Conditioned on the timestep $t$, structured object distribution instruction (SODI) $c_{t}$, and iso-spacing instance map (ISIM) $c_{f}$, the model learns a denoising network $\epsilon_{\theta}$ to estimate the injected noise in $z_{t}$. This process is formalized as:
$\mathcal{L}=\mathbb{E}_{z_{0},t,c_{t},c_{f},\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,c_{t},c_{f})\right\|_{2}^{2}\right]$ (2)
where $\mathcal{L}$ denotes the loss function for training the diffusion model.
@@ IV-B Ablation Study
Effectiveness of SODI. Considering the hallucination problem in deep generative models (DGMs) can cause remote sensing (RS) objects to be inaccurately represented in the background, resulting in inconsistencies between instance-level labels (ISIM) and the generated RS images (as shown in Fig. 6). When low-quality ISIM is used as a control condition for generating images, it leads to a mismatch between the generated images and the provided SODI. This mismatch negatively affects downstream tasks, such as object detection, since the labels for the generated detection data are derived from the SODI. If the generated images do not align with the SODI, the resulting data becomes unreliable, which can significantly impair the model’s performance during training. This discrepancy adversely affects downstream tasks. Therefore, SODI is implemented to guide the image generation process by incorporating a global perspective on image style and instance characteristics, thereby ensuring precision and control. To verify the effectiveness of SODI in ensuring the alignment between ISIM and the generated RS images, we show the testing quantitative metrics of our model with different text conditions in Table I, where ‘NULL’ and ‘BLIP’ denote the text condition is null and an image description generated by BLIP [55]. It can be found that MMO-IG with ‘NULL’ text condition makes it hard to control the accuracy of object class and number on the generated RS images, where the $\rm Acc_{c}$ and $\rm Acc_{n}$ are 92.8 and 93.9 respectively. It means that ISIM, the instance-level label, enjoys low-quality correspondence with the generated RS images, further influencing the following downstream task. $\rm Acc_{c}$ evaluates category alignment between generated instances and SODI-defined classes, while $\rm Acc_{n}$ measures numerical consistency in instance counts relative to SODI specifications.
@@
SCDKG consists of $\rm p_{id}$ matrix that models interdependencies among all objects of different classes (the $\rm p_{id}$ matrix of DIOR dataset is shown in Fig. 7) and $\mathcal{P}_{\rm sgc}$ that formulates diverse spatial geometric characteristics of each object class individually. Notably, $\mathcal{P}_{\rm sgc}$ models the aspect ratio, scale, and location characteristics of RS objects separately. It effectively avoids influence brought by irrational characteristics (e.g., aspect ratio and positional distribution that do not conform to real-world object spatial geometric characteristics visualized in Fig. 8). To verify the effectiveness of SCDKG, we test the performance enhancement brought by each factor in SCDKG and visualize the results in Table II. For $\mathcal{P}_{\rm sgc}$, It can be found that all of the factors can bring performance gains for the generated RS images. Specifically, the modeling of the complex interdependencies among different RS objects makes a more rational object distribution, which helps our model achieve 3.8 and 1.1 improvements in $\rm FID_{zs}$ and CAS on DIOR dataset and 6.8 and 1.0 gains in $\rm FID_{zs}$ and CAS on DIOR-R dataset. Notably, $\rm FID_{zs}$ and $\rm CAS_{zs}$ denote the zero-shot FID and zero-shot CAS, which is introduced to evaluate the quality of the generated images that not in the test set. The improvements brought by $\rm p_{id}$ matrix demonstrate the effectiveness of the interdependencies modeling.
@@
Besides, the rational characteristics of the object aspect ratio, scale, and location bring 8.8, 5.8, and 14.0 improvements in $\rm FID_{zs}$ and 1.4, 3.1, and 5.6 gains in $\rm CAS_{zs}$. This is mainly because the synthesized images enjoy more rational and realistic object spatial characteristics than real RS images, which makes it easier for our model to learn the feature distribution and bring significant performance improvements. Meanwhile, the combination of the three characteristics also can bring further improvements in both $\rm FID_{zs}$ and $\rm CAS_{zs}$. Specifically, the model equipped with SCDKG gains 26.7 and 17.6 in $\rm FID_{zs}$ and $\rm CAS_{zs}$ on DIOR dataset and 29.4 and 19.5 in $\rm FID_{zs}$ and $\rm CAS_{zs}$ on DIOR-R dataset. Significant performance improvements verify the effectiveness of the proposed SCDKG in the generation task of RS images containing MMOs. Some qualified samples are shown in Fig. 9, it can be found that SCDKG enables ensuring rational spatial geometric characteristics for objects with different classes to enhance the reality of generated images.